Question 7 (5 points) Let f(x) = 24 and -2x, x < 5 9 3 22, x > 5 Evaluate(gof)(7) A/
Answer #1

Solutiongi Criven ll 24 (at) -28 ja) x 252 x715 COgof(7) [f(7) gof(7) = = g(34) = 3 (4) a fof(7) from regionis = 9 (4) Hese 224 lie x 25 gof(2) = -2X4 -8 Ang gof (7)

answered by: ANURANJAN SARSAM
Answer #2

Solutiongi Criven ll 24 (at) -28 ja) x 252 x715 COgof(7) [f(7) gof(7) = = g(34) = 3 (4) a fof(7) from regionis = 9 (4) Hese 224 lie x 25 gof(2) = -2X4 -8 Ang gof (7)